Find: (4/9) ÷ 5 Find: (4/9) ÷ 5 Find: (4/9) ÷ 5 Find: (4/9) ÷ 5

Find:
 (4/9) ÷ 5
Find:
 (4/9) ÷ 5
Find:
 (4/9) ÷ 5
Find:
 (4/9) ÷ 5

Grade:12

1 Answers

Harshit Singh
2 years ago
Dear Student

We have,
= (4/9) × reciprocal of 5
= (4/9) × (1/5)
= (4 × 1) / (9 × 5)
= 4/45

Thanks